# Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis

### v3

@@ Appendix G Additional Robustness Result
Figure 11: Comparison of average EM scores across models under varying few-shot settings (0-shot, 1-shot, 2-shot, and 3-shot).
Figure 12: The average cumulative EM score across three different datasets for the different few shot comparisons for various perturbations with three different datasets.
The table 1 summarizes the impact of various perturbation operations on the accuracy and robustness of larger and smaller LLMs models when performing TQA across three distinct datasets (WTQ, TAT-QA, and SCITAB). For each perturbation category, the EM scores typically drop compared to the original configuration, as indicated by negative EmD values. This performance dip is especially notable on the WTQ dataset, where even minor perturbations lead to reduced EM, and certain operations like column transpose and NT conditions cause pronounced declines in EM. Meanwhile, the VP metrics show that predictions shift more substantially under these transformations, signaling sensitivity in both large and small models to altered table structures. Although large models usually have higher baseline EM, they still exhibit performance degradation under perturbations. Particularly striking are conditions like NT, where the table is effectively removed, resulting in EM plummeting to near-zero and high VP, underscoring the models’ heavy reliance on the given tabular data. The results highlight the need for improved architectural and training strategies aimed at enhancing structural awareness and mitigating the negative effects of domain-specific perturbations, thereby making LLMs more robust and reliable in real-world table comprehension scenarios.
Figure 13: The average EM for different models across the TAT-QA dataset with 3-fewshot settings. The models are categorized by their types (Llama2, Llama3, Mistral) and separated by their sizes (7b, 70b, etc.)
Figure 14: Average EM scores for both original and instruction or conversation fine-tuned models on the SCITAB dataset (3-shot setting) under various table augmentation techniques.
G.1 Does newer models have better TQA abilities?
Fig. 15 shows that Llama3 models generally outperform the Llama2 and Mistral models across different configurations, indicating that newer architectures like Llama3 are more effective at table reasoning tasks. The 70b versions of these models generally perform better than their 7b variants, indicating that larger model sizes enhance reasoning capabilities. Overall, the advancements in model architecture and increased model size significantly contribute to better TQA abilities.
Figure 15: The average EM for different models across the WTQ dataset with 3-fewshot settings. The models are categorized by their types (Llama2, Llama3, Mistral) and separated by their sizes (7b, 70b, etc.)
How does model size affect TQA? Larger models (e.g., Llama3-70b, Mixtral-8x7b) generally show higher performance than smaller models (e.g., Llama2-7b, Mistral-7b), as seen in both the bar plot (Fig 15) and heat maps (Fig 10). For instance, Llama3-70b and Mixtral-8x7b have higher EM scores than Llama2-7b and Mistral-7b. This indicates that model size contributes significantly to TQA performance.
Does instruction tuning help TQA? The heat maps in Fig. 10 illustrate the performance of various LLMs, demonstrating that models fine-tuned for instructions or conversation exhibit improved performance. For instance, the Llama3-70b-Instruct model significantly outperforms its original version across all table transformations, indicating that instruction-based fine-tuning enhances the model’s ability to handle complex reasoning tasks. Similarly, conversation-focused fine-tuning also leads to better scores, though with a less pronounced improvement compared to instruction-focused tuning. This suggests that fine-tuning models on specific tasks like following instructions or conversing effectively enhances their capability to interpret and manipulate tabular data, making such approaches valuable for improving performance in structured data tasks. Figure 10(right) distinctly indicates that LLMs that have undergone instruction or conversation-based fine-tuning outperform their base counterparts in TQA tasks for SCITAB dataset. Although SCITAB is inherently a classification task presented in a TQA format, using few-shot prompting provides valuable context, ultimately leading to more accurate and relevant responses.
Appendix H Attention Matrix Measure
Sink rate: Gu et al. introduced the metric that measures the proportion of attention that “collapses" onto a specific token or a small set of tokens. A high sink rate indicates that the attention head focuses very narrowly, potentially ignoring other parts of the input.
$\text{Sink Rate}_{\epsilon}=\frac{1}{T}\sum_{j}\alpha_{1,j}^{(\ell,h)}>\epsilon,$
where $\alpha_{1,j}^{(\ell,h)}$ is the attention weight from the first token (token with index 1(<bos>)) to the j-th token in the $\ell$-th layer and $h$-th attention head.
Entropy: In the context of attention, entropy measures the diversity or distribution of attention across all tokens. High entropy means the attention is distributed more evenly, while low entropy suggests a concentrated attention pattern.
$H_{i}=-\sum_{j}\textbf{A}_{ij}\log(\textbf{A}_{ij})$
where $\textbf{A}_{ij}$ is the attention weight to the $j$-th token.
Frobenius Norm: The Frobenius norm of the attention matrix provides a measure of the overall magnitude or “strength” of attention. It is calculated as the square root of the sum of the squares of all elements in the matrix, indicating the total amount of attention being distributed.
$||\textbf{A}||_{F}=\sqrt{\sum_{i}\sum_{j}\textbf{A}_{ij}^{2}}$
Norm Difference: The difference between the attention matrices computed on the original (dev) and perturbed table(pert):
$\mathrm{NormDiff}\;=\;\bigl{\|}\mathbf{A}^{\text{dev}}\;-\;\mathbf{A}^{\text{pert}}\bigr{\|}_{F},$
where $\mathbf{A}^{dev}$ and $\mathbf{A}^{pert}$ are the attention matrices tokenized using the original table (‘dev”) and the perturbed table, respectively, and $\lVert\cdot\rVert_{F}$ is the Frobenius norm.
H.1 Cross-Model and Dataset Analysis of Perturbation Sensitivity
Figure 16: Spearman correlation between perturbation-induced attention-entropy change and EM degradation across 18 model–dataset pairs.
[…]

### v4

@@ Appendix G Additional Robustness Result
Table 4: EM results across all datasets (WTQ, TAT-QA, and SCITAB) and model families. Note: The labels B, C, and I refer to the Base, Chat, and Instruct variants of the model.
Model Original Column Swap Row Swap Transpose Transpose Row Swap Transpose Col Swap NT DVP RVP NVP
WTQ
Llama2 7B 0.15 ± 0.01 0.148 ± 0.01 0.118 ± 0.018 0.11 ± 0.012 0.108 ± 0.015 0.078 ± 0.022 0.021 ± 0.003 0.033 ± 0.004 0.012 ± 0.008 0.014 ± 0.007
Llama3 8B 0.27 ± 0.035 0.225 ± 0.008 0.221 ± 0.0 0.216 ± 0.0 0.181 ± 0.015 0.167 ± 0.014 0.016 ± 0.0 0.149 ± 0.03 0.085 ± 0.0 0.021 ± 0.0
Mistral 7B 0.294 ± 0.0 0.291 ± 0.006 0.224 ± 0.006 0.234 ± 0.015 0.214 ± 0.003 0.194 ± 0.007 0.025 ± 0.0 0.118 ± 0.004 0.066 ± 0.018 0.031 ± 0.004
Qwen2.5 7B 0.191 ± 0.013 0.196 ± 0.025 0.155 ± 0.012 0.158 ± 0.003 0.16 ± 0.006 0.131 ± 0.02 0.023 ± 0.011 0.064 ± 0.014 0.009 ± 0.004 0.026 ± 0.022
Qwen3 8B B 0.297 ± 0.02 0.246 ± 0.03 0.187 ± 0.02 0.272 ± 0.015 0.239 ± 0.012 0.168 ± 0.012 0.026 ± 0.016 0.139 ± 0.018 0.057 ± 0.025 0.031 ± 0.004
Llama2 70B 0.273 ± 0.007 0.299 ± 0.013 0.26 ± 0.015 0.235 ± 0.032 0.229 ± 0.01 0.178 ± 0.007 0.044 ± 0.013 0.135 ± 0.021 0.061 ± 0.004 0.038 ± 0.008
Llama3 70B 0.368 ± 0.031 0.36 ± 0.04 0.311 ± 0.032 0.348 ± 0.0 0.319 ± 0.021 0.25 ± 0.04 0.082 ± 0.01 0.199 ± 0.021 0.125 ± 0.027 0.085 ± 0.012
Mistral 8x7B 0.281 ± 0.02 0.291 ± 0.012 0.242 ± 0.01 0.265 ± 0.013 0.255 ± 0.008 0.219 ± 0.011 0.034 ± 0.0 0.147 ± 0.004 0.073 ± 0.004 0.05 ± 0.0
Qwen2.5 72B 0.37 ± 0.021 0.369 ± 0.03 0.306 ± 0.019 0.366 ± 0.041 0.375 ± 0.016 0.235 ± 0.013 0.02 ± 0.0 0.227 ± 0.012 0.137 ± 0.015 0.043 ± 0.007
Qwen3 30B B 0.265 ± 0.01 0.237 ± 0.025 0.225 ± 0.0 0.252 ± 0.014 0.209 ± 0.014 0.163 ± 0.011 0.015 ± 0.008 0.13 ± 0.02 0.076 ± 0.011 0.035 ± 0.007
Llama2 7B C 0.225 ± 0.014 0.233 ± 0.003 0.176 ± 0.014 0.181 ± 0.0 0.142 ± 0.007 0.123 ± 0.007 0.029 ± 0.006 0.121 ± 0.0 0.064 ± 0.01 0.011 ± 0.005
Llama3 8B I 0.373 ± 0.013 0.348 ± 0.005 0.33 ± 0.019 0.315 ± 0.02 0.314 ± 0.013 0.242 ± 0.019 0.01 ± 0.01 0.253 ± 0.022 0.168 ± 0.011 0.045 ± 0.008
Mistral 7B I 0.325 ± 0.022 0.324 ± 0.0 0.248 ± 0.007 0.284 ± 0.018 0.275 ± 0.005 0.189 ± 0.01 0.015 ± 0.0 0.248 ± 0.007 0.229 ± 0.008 0.019 ± 0.004
Qwen2.5 7B I 0.335 ± 0.012 0.312 ± 0.01 0.292 ± 0.02 0.273 ± 0.025 0.242 ± 0.012 0.19 ± 0.012 0.002 ± 0.003 0.236 ± 0.03 0.158 ± 0.015 0.057 ± 0.007
Qwen3 8B 0.295 ± 0.017 0.246 ± 0.03 0.187 ± 0.02 0.272 ± 0.015 0.239 ± 0.012 0.168 ± 0.012 0.01 ± 0.013 0.139 ± 0.018 0.057 ± 0.025 0.031 ± 0.004
Llama2 70B C 0.306 ± 0.02 0.302 ± 0.017 0.248 ± 0.01 0.25 ± 0.008 0.253 ± 0.011 0.201 ± 0.018 0.03 ± 0.0 0.187 ± 0.022 0.109 ± 0.027 0.045 ± 0.004
Llama3 70B I 0.544 ± 0.01 0.533 ± 0.006 0.447 ± 0.015 0.528 ± 0.016 0.497 ± 0.012 0.387 ± 0.013 0.083 ± 0.005 0.407 ± 0.015 0.314 ± 0.018 0.095 ± 0.011
Mistral 8x7B I 0.338 ± 0.015 0.317 ± 0.006 0.273 ± 0.015 0.327 ± 0.022 0.279 ± 0.005 0.225 ± 0.013 0.049 ± 0.0 0.149 ± 0.007 0.097 ± 0.004 0.061 ± 0.004
Qwen2.5 72B I 0.499 ± 0.016 0.501 ± 0.023 0.422 ± 0.013 0.472 ± 0.02 0.487 ± 0.008 0.365 ± 0.016 0.062 ± 0.003 0.369 ± 0.007 0.191 ± 0.007 0.061 ± 0.008
Qwen3 30B 0.261 ± 0.012 0.225 ± 0.022 0.204 ± 0.03 0.24 ± 0.027 0.209 ± 0.01 0.173 ± 0.003 0.018 ± 0.019 0.137 ± 0.022 0.083 ± 0.015 0.038 ± 0.004
TAPAS Tiny (WTQ) 0.562 0.507 0.230 0.078 0.064 0.078 0.355 0.418 0.064
TAPAS Small (WTQ) 0.700 0.691 0.389 0.078 0.083 0.049 0.645 0.603 0.064
TAPAS Mini (WTQ) 0.700 0.666 0.398 0.113 0.078 0.059 0.596 0.610 0.071
TAPAS Medium (WTQ) 0.720 0.686 0.478 0.074 0.059 0.069 0.695 0.624 0.057
TAPAS Base (WTQ) 0.705 0.696 0.465 0.083 0.044 0.088 0.695 0.603 0.050
TAPAS Large (WTQ) 0.713 0.700 0.485 0.116 0.096 0.103 0.702 0.596 0.057
TAPEX Base (WTQ) 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
TAPEX Large (WTQ) 0.098 0.078 0.079 0.020 0.034 0.025 0.078 0.057 0.000
TATQ-A
Llama2 7B 0.184 ± 0.007 0.16 ± 0.008 0.154 ± 0.005 0.147 ± 0.001 0.119 ± 0.007 0.133 ± 0.001 0.004 ± 0.001
Llama3 8B 0.298 ± 0.002 0.248 ± 0.005 0.26 ± 0.004 0.279 ± 0.006 0.218 ± 0.014 0.252 ± 0.006 0.003 ± 0.0
Mistral 7B 0.27 ± 0.003 0.237 ± 0.001 0.25 ± 0.005 0.239 ± 0.004 0.186 ± 0.002 0.216 ± 0.003 0.008 ± 0.0
Qwen2.5 7B 0.226 ± 0.004 0.195 ± 0.006 0.197 ± 0.007 0.216 ± 0.005 0.181 ± 0.013 0.208 ± 0.004 0.003 ± 0.003
Qwen3 8B B 0.254 ± 0.005 0.229 ± 0.017 0.218 ± 0.002 0.236 ± 0.007 0.208 ± 0.012 0.216 ± 0.009 0.0 ± 0.0
Llama2 70B 0.314 ± 0.016 0.283 ± 0.022 0.275 ± 0.007 0.263 ± 0.01 0.222 ± 0.019 0.267 ± 0.008 0.005 ± 0.003
Llama3 70B 0.384 ± 0.011 0.36 ± 0.012 0.362 ± 0.015 0.357 ± 0.016 0.322 ± 0.004 0.339 ± 0.01 0.014 ± 0.001
Mistral 8x7B 0.317 ± 0.015 0.279 ± 0.002 0.273 ± 0.003 0.284 ± 0.01 0.215 ± 0.011 0.275 ± 0.004 0.01 ± 0.0
Qwen2.5 72B 0.33 ± 0.002 0.3 ± 0.008 0.306 ± 0.005 0.305 ± 0.003 0.268 ± 0.024 0.307 ± 0.012 0.003 ± 0.003
Qwen3 30B B 0.336 ± 0.011 0.302 ± 0.019 0.304 ± 0.004 0.32 ± 0.026 0.239 ± 0.008 0.28 ± 0.008 0.006 ± 0.003
Llama2 7B C 0.185 ± 0.004 0.159 ± 0.003 0.168 ± 0.007 0.155 ± 0.005 0.124 ± 0.003 0.135 ± 0.003 0.008 ± 0.004
Llama3 8B I 0.278 ± 0.004 0.238 ± 0.002 0.253 ± 0.001 0.268 ± 0.006 0.215 ± 0.007 0.247 ± 0.0 0.001 ± 0.001
Mistral 7B I 0.229 ± 0.006 0.188 ± 0.005 0.197 ± 0.004 0.208 ± 0.002 0.17 ± 0.003 0.191 ± 0.002 0.001 ± 0.0
Qwen2.5 7B I 0.24 ± 0.007 0.222 ± 0.017 0.219 ± 0.005 0.227 ± 0.003 0.197 ± 0.011 0.218 ± 0.009 0.0 ± 0.0
Qwen3 8B 0.253 ± 0.006 0.241 ± 0.009 0.257 ± 0.006 0.249 ± 0.008 0.216 ± 0.001 0.226 ± 0.018 0.007 ± 0.004
Llama2 70B C 0.25 ± 0.01 0.211 ± 0.008 0.226 ± 0.004 0.211 ± 0.006 0.172 ± 0.003 0.197 ± 0.012 0.002 ± 0.002
Llama3 70B I 0.441 ± 0.004 0.417 ± 0.006 0.405 ± 0.005 0.431 ± 0.001 0.37 ± 0.006 0.391 ± 0.005 0.014 ± 0.001
Mistral 8x7B I 0.317 ± 0.008 0.283 ± 0.017 0.273 ± 0.004 0.27 ± 0.006 0.22 ± 0.01 0.271 ± 0.013 0.013 ± 0.0
Qwen2.5 72B I 0.353 ± 0.009 0.353 ± 0.005 0.337 ± 0.006 0.34 ± 0.012 0.303 ± 0.006 0.33 ± 0.007 0.008 ± 0.0
Qwen3 30B 0.346 ± 0.022 0.305 ± 0.013 0.304 ± 0.027 0.298 ± 0.018 0.258 ± 0.008 0.26 ± 0.016 0.01 ± 0.0
TAPAS Tiny (WTQ) 0.017 0.012 0.010 0.040 0.025 0.035
TAPAS Small (WTQ) 0.032 0.030 0.022 0.093 0.052 0.092
TAPAS Mini (WTQ) 0.045 0.046 0.015 0.105 0.057 0.097
TAPAS Medium (WTQ) 0.046 0.046 0.019 0.125 0.075 0.132
TAPAS Base (WTQ) 0.039 0.040 0.021 0.071 0.054 0.065
TAPAS Large (WTQ) 0.037 0.045 0.029 0.085 0.058 0.087
TAPEX Base (WTQ) 0.010 0.020 0.008 0.012 0.012 0.016
TAPEX Large (WTQ) 0.012 0.008 0.005 0.002 0.008 0.009
SCITAB
Llama2 7B 0.001 ± 0.001 0.003 ± 0.004 0.002 ± 0.0 0.002 ± 0.001 0.002 ± 0.003 0.001 ± 0.001 0.0 ± 0.0
Llama3 8B 0.338 ± 0.025 0.349 ± 0.03 0.32 ± 0.016 0.342 ± 0.011 0.352 ± 0.03 0.319 ± 0.017 0.0 ± 0.0
Mistral 7B 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0
Qwen2.5 7B 0.002 ± 0.004 0.0 ± 0.0 0.002 ± 0.001 0.002 ± 0.003 0.003 ± 0.004 0.002 ± 0.001 0.0 ± 0.0
Qwen3 8B B 0.327 ± 0.036 0.293 ± 0.013 0.301 ± 0.019 0.312 ± 0.011 0.28 ± 0.007 0.278 ± 0.017 0.0 ± 0.0
Llama2 70B 0.034 ± 0.007 0.027 ± 0.028 0.022 ± 0.02 0.04 ± 0.009 0.036 ± 0.013 0.031 ± 0.008 0.0 ± 0.0
Llama3 70B 0.38 ± 0.03 0.366 ± 0.007 0.35 ± 0.009 0.367 ± 0.007 0.364 ± 0.007 0.356 ± 0.043 0.0 ± 0.0
Mistral 8x7B 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0 0.0 ± 0.0
Qwen2.5 72B 0.349 ± 0.019 0.3 ± 0.035 0.293 ± 0.045 0.341 ± 0.036 0.311 ± 0.005 0.359 ± 0.058 0.0 ± 0.0
Qwen3 30B B 0.093 ± 0.012 0.065 ± 0.011 0.074 ± 0.01 0.082 ± 0.009 0.078 ± 0.007 0.059 ± 0.01 0.0 ± 0.0
Llama2 7B C 0.384 ± 0.006 0.387 ± 0.008 0.382 ± 0.0 0.378 ± 0.005 0.383 ± 0.008 0.387 ± 0.004 0.0 ± 0.0
Llama3 8B I 0.32 ± 0.003 0.313 ± 0.017 0.322 ± 0.008 0.315 ± 0.014 0.324 ± 0.004 0.311 ± 0.009 0.0 ± 0.0
Mistral 7B I 0.284 ± 0.012 0.28 ± 0.016 0.28 ± 0.013 0.282 ± 0.018 0.279 ± 0.009 0.283 ± 0.011 0.0 ± 0.0
Qwen2.5 7B I 0.352 ± 0.003 0.322 ± 0.007 0.353 ± 0.013 0.364 ± 0.01 0.347 ± 0.004 0.339 ± 0.004 0.0 ± 0.0
Qwen3 8B 0.378 ± 0.016 0.341 ± 0.013 0.325 ± 0.011 0.378 ± 0.034 0.356 ± 0.005 0.322 ± 0.009 0.0 ± 0.0
Llama2 70B C 0.084 ± 0.004 0.072 ± 0.01 0.076 ± 0.005 0.091 ± 0.013 0.077 ± 0.008 0.078 ± 0.014 0.0 ± 0.0
Llama3 70B I 0.475 ± 0.007 0.462 ± 0.014 0.457 ± 0.006 0.461 ± 0.014 0.454 ± 0.017 0.451 ± 0.014 0.0 ± 0.0
Mistral 8x7B I 0.083 ± 0.001 0.078 ± 0.008 0.08 ± 0.004 0.084 ± 0.007 0.087 ± 0.008 0.082 ± 0.003 0.0 ± 0.0
Qwen2.5 72B I 0.63 ± 0.012 0.578 ± 0.011 0.579 ± 0.01 0.626 ± 0.023 0.575 ± 0.011 0.6 ± 0.014 0.0 ± 0.0
Qwen3 30B 0.021 ± 0.008 0.022 ± 0.012 0.012 ± 0.001 0.008 ± 0.0 0.015 ± 0.002 0.007 ± 0.004 0.0 ± 0.0
Across all three datasets—WTQ, TAT-QA, and SCITAB—the EM results in Table 4 elaborate on the consistent trends supporting the main findings. In addition to standard deviation for statistical significance, it also includes the TAPAS models(ranging from Tiny to Large) that are fine-tuned on the WTQ dataset and TAPEX models(Large and Base) fine-tuned on WTQ dataset. The TAPAS models exhibit a distinctive performance profile that contrasts sharply with the general-purpose LLMs evaluated in the main text. On WTQ, TAPAS achieves notably high EM scores on the original setting (0.562–0.720) and retains relatively strong performance under most structural perturbations, with only modest declines for operations such as column swap and row swap. However, its robustness sharply deteriorates under value-based perturbations such as DVP, RVP, and NVP, with EM drops similar to those seen for many instruction-tuned LLMs. On TAT-QA, TAPAS’s performance is substantially lower than on WTQ, with EM scores often under 0.05 for most structural perturbations—an order-of-magnitude drop that reflects limited domain transferability from its trained dataset. Overall, TAPAS models show a better structural robustness within their training domain (WTQ) but exhibit steep cross-domain degradation and heightened vulnerability to semantic content changes, reinforcing the main-text conclusion that while specialized tabular models can excel in-domain, instruction-tuned LLMs demonstrate superior adaptability and broader robustness across perturbations and domains. Across all datasets, TAPEX performs substantially worse than both TAPAS and modern LLM-based models, with EM scores often near zero regardless of perturbation. TAPEX models — even the WTQ-specific large variant, peak at only 0.098 EM and collapse under perturbations, indicating a lack of generalization, likely due to training, evaluation mismatches, or overfitting. On TAT-QA, LLM-based models again lead (e.g., Llama3-70B I at 0.441), whereas TAPAS drops to the 0.017–0.046 range and TAPEX remains near zero.
G.1 Score with F1 Measure
Table 5: F1 measure results across all datasets (WTQ, TAT-QA, and SCITAB) and model families. Note: The labels B, C, and I refer to the Base, Chat, and Instruct variants of the model.
Model Original Column Swap Row Swap Transpose Transpose Row Swap Transpose Col Swap NT DVP RVP NVP
WTQ
Llama2 7B 0.173 ± 0.008 0.169 ± 0.01 0.134 ± 0.014 0.127 ± 0.009 0.129 ± 0.014 0.09 ± 0.021 0.032 ± 0.004 0.053 ± 0.01 0.022 ± 0.012 0.028 ± 0.007
Llama3 8B 0.316 ± 0.021 0.254 ± 0.012 0.252 ± 0.0 0.259 ± 0.0 0.212 ± 0.02 0.193 ± 0.014 0.027 ± 0.002 0.187 ± 0.035 0.112 ± 0.0 0.027 ± 0.0
Mistral 7B 0.322 ± 0.008 0.321 ± 0.008 0.252 ± 0.004 0.262 ± 0.014 0.243 ± 0.003 0.214 ± 0.006 0.037 ± 0.005 0.139 ± 0.004 0.083 ± 0.019 0.036 ± 0.002
Qwen2.5 7B 0.271 ± 0.015 0.254 ± 0.017 0.206 ± 0.014 0.206 ± 0.005 0.212 ± 0.006 0.161 ± 0.019 0.042 ± 0.011 0.111 ± 0.011 0.044 ± 0.004 0.04 ± 0.023
Qwen3 8B B 0.35 ± 0.01 0.302 ± 0.03 0.252 ± 0.021 0.318 ± 0.021 0.286 ± 0.015 0.191 ± 0.013 0.043 ± 0.022 0.173 ± 0.016 0.091 ± 0.03 0.039 ± 0.005
Llama2 70B 0.308 ± 0.006 0.331 ± 0.013 0.292 ± 0.011 0.268 ± 0.032 0.246 ± 0.011 0.2 ± 0.007 0.038 ± 0.009 0.15 ± 0.023 0.079 ± 0.007 0.054 ± 0.014
Llama3 70B 0.414 ± 0.036 0.4 ± 0.033 0.354 ± 0.03 0.38 ± 0.008 0.352 ± 0.021 0.277 ± 0.038 0.105 ± 0.01 0.217 ± 0.026 0.139 ± 0.022 0.109 ± 0.01
Mistral 8x7B 0.305 ± 0.023 0.313 ± 0.012 0.258 ± 0.004 0.291 ± 0.012 0.275 ± 0.004 0.239 ± 0.009 0.04 ± 0.001 0.162 ± 0.006 0.082 ± 0.005 0.061 ± 0.005
Qwen2.5 72B 0.432 ± 0.019 0.426 ± 0.024 0.364 ± 0.018 0.415 ± 0.04 0.431 ± 0.011 0.27 ± 0.008 0.037 ± 0.002 0.25 ± 0.015 0.159 ± 0.02 0.069 ± 0.007
Qwen3 30B B 0.313 ± 0.016 0.272 ± 0.013 0.275 ± 0.017 0.299 ± 0.014 0.247 ± 0.018 0.202 ± 0.01 0.045 ± 0.01 0.168 ± 0.025 0.108 ± 0.005 0.059 ± 0.009
Llama2 7B C 0.269 ± 0.008 0.271 ± 0.008 0.204 ± 0.022 0.209 ± 0.004 0.171 ± 0.005 0.145 ± 0.01 0.057 ± 0.003 0.146 ± 0.0 0.078 ± 0.02 0.04 ± 0.006
Llama3 8B I 0.46 ± 0.014 0.434 ± 0.008 0.41 ± 0.027 0.386 ± 0.016 0.368 ± 0.007 0.292 ± 0.016 0.022 ± 0.016 0.336 ± 0.023 0.235 ± 0.004 0.072 ± 0.005
Mistral 7B I 0.417 ± 0.016 0.409 ± 0.004 0.343 ± 0.01 0.371 ± 0.016 0.351 ± 0.006 0.257 ± 0.016 0.025 ± 0.0 0.327 ± 0.004 0.289 ± 0.012 0.065 ± 0.003
Qwen2.5 7B I 0.459 ± 0.007 0.442 ± 0.009 0.397 ± 0.036 0.415 ± 0.023 0.366 ± 0.02 0.306 ± 0.011 0.026 ± 0.003 0.335 ± 0.026 0.233 ± 0.006 0.09 ± 0.011
Qwen3 8B 0.347 ± 0.011 0.302 ± 0.03 0.252 ± 0.021 0.318 ± 0.021 0.286 ± 0.015 0.191 ± 0.013 0.032 ± 0.014 0.173 ± 0.016 0.091 ± 0.03 0.039 ± 0.005
Llama2 70B C 0.339 ± 0.011 0.341 ± 0.01 0.282 ± 0.012 0.272 ± 0.005 0.281 ± 0.008 0.215 ± 0.033 0.024 ± 0.0 0.214 ± 0.02 0.134 ± 0.036 0.056 ± 0.007
Llama3 70B I 0.619 ± 0.006 0.605 ± 0.014 0.531 ± 0.007 0.611 ± 0.014 0.574 ± 0.024 0.466 ± 0.013 0.109 ± 0.007 0.491 ± 0.005 0.383 ± 0.018 0.13 ± 0.011
Mistral 8x7B I 0.406 ± 0.006 0.38 ± 0.008 0.319 ± 0.02 0.386 ± 0.02 0.334 ± 0.005 0.253 ± 0.012 0.057 ± 0.0 0.203 ± 0.014 0.113 ± 0.004 0.072 ± 0.008
Qwen2.5 72B I 0.589 ± 0.013 0.59 ± 0.01 0.511 ± 0.005 0.551 ± 0.012 0.563 ± 0.019 0.432 ± 0.004 0.091 ± 0.005 0.475 ± 0.004 0.265 ± 0.01 0.103 ± 0.011
Qwen3 30B 0.368 ± 0.025 0.345 ± 0.025 0.301 ± 0.029 0.34 ± 0.027 0.304 ± 0.008 0.245 ± 0.021 0.041 ± 0.017 0.24 ± 0.007 0.187 ± 0.01 0.064 ± 0.01
TAPAS Tiny (WTQ) 0.584 0.525 0.282 0.094 0.079 0.091 0.390 0.451 0.064
TAPAS Small (WTQ) 0.717 0.705 0.422 0.087 0.098 0.061 0.665 0.621 0.069
TAPAS Mini (WTQ) 0.717 0.683 0.423 0.120 0.087 0.074 0.635 0.633 0.082
TAPAS Medium (WTQ) 0.737 0.704 0.498 0.085 0.062 0.078 0.712 0.643 0.064
TAPAS Base (WTQ) 0.720 0.710 0.484 0.097 0.060 0.105 0.706 0.624 0.076
TAPAS Large (WTQ) 0.735 0.716 0.519 0.120 0.096 0.109 0.709 0.611 0.089
TAPEX Base (WTQ) 0.020 0.025 0.039 0.012 0.014 0.029 0.022 0.061 0.004
TAPEX Large (WTQ) 0.120 0.093 0.100 0.032 0.048 0.040 0.089 0.068 0.018
TATQ-A
Llama2 7B 0.216 ± 0.007 0.191 ± 0.01 0.185 ± 0.004 0.179 ± 0.001 0.146 ± 0.005 0.163 ± 0.004 0.014 ± 0.001
Llama3 8B 0.348 ± 0.003 0.294 ± 0.004 0.306 ± 0.002 0.326 ± 0.005 0.264 ± 0.014 0.297 ± 0.006 0.015 ± 0.002
Mistral 7B 0.329 ± 0.002 0.292 ± 0.001 0.306 ± 0.007 0.295 ± 0.003 0.235 ± 0.004 0.272 ± 0.004 0.021 ± 0.0
Qwen2.5 7B 0.282 ± 0.002 0.254 ± 0.009 0.251 ± 0.005 0.27 ± 0.008 0.233 ± 0.014 0.264 ± 0.003 0.02 ± 0.002
Qwen3 8B B 0.304 ± 0.005 0.28 ± 0.016 0.27 ± 0.005 0.285 ± 0.01 0.256 ± 0.011 0.266 ± 0.009 0.022 ± 0.001
Llama2 70B 0.378 ± 0.015 0.349 ± 0.022 0.34 ± 0.006 0.33 ± 0.01 0.283 ± 0.02 0.328 ± 0.011 0.013 ± 0.004
Llama3 70B 0.45 ± 0.011 0.423 ± 0.013 0.424 ± 0.01 0.421 ± 0.018 0.389 ± 0.002 0.403 ± 0.011 0.054 ± 0.0
Mistral 8x7B 0.364 ± 0.014 0.322 ± 0.004 0.319 ± 0.001 0.333 ± 0.01 0.256 ± 0.013 0.325 ± 0.003 0.031 ± 0.0
Qwen2.5 72B 0.388 ± 0.002 0.358 ± 0.002 0.361 ± 0.006 0.365 ± 0.008 0.327 ± 0.02 0.364 ± 0.018 0.02 ± 0.007
Qwen3 30B B 0.384 ± 0.014 0.351 ± 0.021 0.355 ± 0.004 0.371 ± 0.027 0.29 ± 0.008 0.328 ± 0.005 0.031 ± 0.002
Llama2 7B C 0.23 ± 0.006 0.198 ± 0.003 0.21 ± 0.008 0.199 ± 0.004 0.156 ± 0.005 0.175 ± 0.003 0.028 ± 0.012
Llama3 8B I 0.35 ± 0.006 0.305 ± 0.003 0.321 ± 0.0 0.34 ± 0.006 0.279 ± 0.006 0.311 ± 0.002 0.005 ± 0.002
Mistral 7B I 0.312 ± 0.002 0.261 ± 0.003 0.272 ± 0.002 0.286 ± 0.0 0.24 ± 0.002 0.267 ± 0.003 0.024 ± 0.0
Qwen2.5 7B I 0.306 ± 0.007 0.285 ± 0.015 0.281 ± 0.007 0.294 ± 0.003 0.26 ± 0.013 0.284 ± 0.009 0.018 ± 0.002
Qwen3 8B 0.316 ± 0.008 0.301 ± 0.008 0.321 ± 0.005 0.32 ± 0.009 0.279 ± 0.002 0.287 ± 0.017 0.033 ± 0.004
Llama2 70B C 0.307 ± 0.011 0.268 ± 0.009 0.282 ± 0.007 0.27 ± 0.007 0.221 ± 0.004 0.248 ± 0.008 0.004 ± 0.003
Llama3 70B I 0.523 ± 0.005 0.499 ± 0.005 0.486 ± 0.005 0.514 ± 0.001 0.452 ± 0.006 0.472 ± 0.005 0.054 ± 0.0
Mistral 8x7B I 0.372 ± 0.015 0.336 ± 0.016 0.327 ± 0.004 0.326 ± 0.007 0.269 ± 0.01 0.325 ± 0.011 0.036 ± 0.0
Qwen2.5 72B I 0.426 ± 0.008 0.426 ± 0.006 0.411 ± 0.009 0.418 ± 0.01 0.377 ± 0.004 0.4 ± 0.005 0.04 ± 0.0
Qwen3 30B 0.409 ± 0.019 0.373 ± 0.02 0.373 ± 0.035 0.371 ± 0.021 0.319 ± 0.01 0.329 ± 0.021 0.034 ± 0.0
TAPAS Tiny (WTQ) 0.031 0.026 0.026 0.047 0.035 0.046
TAPAS Small (WTQ) 0.050 0.043 0.031 0.100 0.060 0.099
TAPAS Mini (WTQ) 0.053 0.052 0.024 0.112 0.062 0.103
TAPAS Medium (WTQ) 0.057 0.061 0.032 0.145 0.084 0.144
TAPAS Base (WTQ) 0.052 0.049 0.033 0.083 0.063 0.081
TAPAS Large (WTQ) 0.045 0.052 0.039 0.106 0.074 0.115
TAPEX Base (WTQ) 0.013 0.021 0.010 0.013 0.013 0.020
TAPEX Large (WTQ) 0.020 0.017 0.012 0.009 0.013 0.016
SCITAB
Llama2 7B 0.001 ± 0.002 0.003 ± 0.004 0.002 ± 0.0 0.002 ± 0.001 0.002 ± 0.003 0.001 ± 0.002 0.0 ± 0.0
Llama3 8B 0.341 ± 0.024 0.351 ± 0.031 0.323 ± 0.016 0.345 ± 0.011 0.354 ± 0.03 0.32 ± 0.018 0.0 ± 0.0
[…]